OpenStudy (anonymous):

Yolanda has finished 45% of the homework problems she was assigned. So far she has finished 27 problems. How many problems was Yolanda assigned? Justify your answer.

OpenStudy (anonymous):

i just mainly need the formula, not an answer

OpenStudy (greatlife44):

\[\frac{ number~of~problems~completed }{ total~problems~ } = \frac{ percent }{ 100 }\]

OpenStudy (anonymous):

so 27/x 45/100

OpenStudy (greatlife44):

:) yes

OpenStudy (anonymous):

and...

OpenStudy (anonymous):

\[45 \% p=27,p=\frac{27}{0.45}=60 \]

OpenStudy (anonymous):

so, i do 27 divided by .45 to get 60?

OpenStudy (anonymous):

Thanks so so so much!

OpenStudy (anonymous):

thanks @greatlife44 too!